A measured 43.2 km/hr equals how many m/s? a. b. 11. 7.30 x 102 m/s 4.00 x 101 m/s c. 12.0 m/s d. 37.0 m/s
Answer #1

11)

1\,{km}=10^3\,m

1\,hr=60\,min=3600\,s

1\,km/hr=\frac{10^3}{3600}\,m/s

43.2\,km/hr=43.2*\frac{10^3}{3600}=12.0\,m/s

Answer is c)\,12.0\,m/s